OpenStudy (anonymous):

HELP FAST I WILL FAN AND MEDAL a bag of fruit contains 4 apples, 1 plum, 2 apricots, and 3 oranges. pieces of fruit are drawn twice with replacement what is p( apple, then apricots)?

OpenStudy (anonymous):

\[A. \frac{ 4 }{ 5 }\] \[B. \frac{ 2 }{25}\] \[C. \frac{ 3 }{ 25 }\] \[D. \frac{ 3 }{ 5}\]

OpenStudy (anonymous):

@123.go @animefreakdesu @beccaboo333 @bloopman @ganeshie8 @iPwnBunnies @Jennaaaaaa @mathmale @Preetha @satellite73 @Shay17 @SolomonZelman @Whitemonsterbunny17 @xGuardians @Zale101

OpenStudy (anonymous):

@xGuardians

OpenStudy (xguardians):

10 Fruits in Total Odds of Apple: 4/10 Odds of Apricot (After Apple): 2/10 4/10 * 2/10 OR (Simplified) 2/5 * 1/5 = 3/25 or 12%

OpenStudy (xguardians):

Sorry misread that it said that there IS replacement :)

OpenStudy (anonymous):

thanxs